# MESA-QUEST: Tracing the formation of direct collapse black hole seeds via quasi-stars

### v2

@@ 6.4 Results
The results of implementing the Dutch wind scheme are plotted in Fig. 3. These winds are now strong enough to severely limit the final black hole masses. We see that the winds are enough to immediately eject the envelope upon expansion in the case of the $\alpha=1.4$ as it is enough to break the equilibrium that is otherwise reached. The $\alpha=1.0$ model lives only marginally longer, shedding its mass, while the black hole rapidly consumes what little remains. Our $\alpha<1.0$ models remain largely unaffected due to both their lower accretion luminosity and their increased temperature. As stated previously, the transition between wind models for the Dutch scheme occurs at $T_{\mathrm{eff}}=10^{4}$ K, which both low $\alpha$ models remain well above for the entirety of their lives. The Dutch scheme then serves as an intermediate case for a wind mass loss rate.
@@
It is of note that envelope accretion can heavily suppress the outflow and sustain the quasi-star’s lifetime, allowing the black hole to accrete more of the envelope before it is shed (C. Dotan et al., 2011; D. Fiacconi & E. M. Rossi, 2015; L. Haemmerlé et al., 2021). This serves as motivation to pursue the addition of an envelope accretion scheme to our models in future work.
7 Discussion & Conclusions
@@
In this work, we have successfully implemented quasi-star models previously created by W. H. Ball (2012) using the Cambridge STARS code (P. P. Eggleton, 1972) into MESA (B. Paxton et al., 2011). This implementation includes the Bondi inner boundary condition as well as the Reimers wind scheme, both of which serve as our fiducial case. We also add the ability to take the accretion rate as a semi-free parameter in order to proxy the addition of other physics. In our fiducial models, we show that final black hole masses of $\sim 10^{3}$ M can be attained even with accretion suppressing effects (e.g., angular momentum).
@@
We have implemented new inner boundary conditions based on a saturated-convection zone around the black hole proposed by E. R. Coughlin & M. C. Begelman (2024). Doing so allows the BH to grow to $\sim$0.55 M, a factor of 5 increase from the previous $\sim$0.11 M achieved using the Bondi sphere as the inner boundary. The quasi-star, under these conditions, then becomes a heavy seed, capable of growing and reaching SMBH masses by early epochs thus explaining recent JWST observations. However, these results change dramatically when accounting for additional physics such as mass loss due to winds.
@@
We have studied three wind schemes in our models: the Reimers, Dutch, and super-Eddington radiation-driven winds. These serve as lower, middle, and upper bounds on mass loss rates for quasi-stars, as the actual values for surface mass loss are unknown without additional hydrodynamic and radiation transport simulations. We also presently do not account for envelope accretion, which is capable of balancing the mass loss and sustaining the envelope for an extended period of time. The Reimers and Dutch schemes may then serve as a proxy for that, given that the radiation-driven wind scheme is likely more realistic and catered to quasi-stars. We find that for the lower bound Reimers wind scheme, little change occurs to the final black hole mass for all accretion factors. The Dutch scheme, however, severely limits the growth of our higher $\alpha$ values, suggesting that if net mass loss occurs in this way that accretion enhancing effects (e.g., magnetic fields and jets) may actually limit the viability of quasi-stars to form heavy seeds. Finally, the radiation-driven wind scheme presents the harshest limit on black hole growth, only allowing for a factor of two growth at most for all rates. If this wind is the most realistic, and accretion onto the envelope is not sufficient enough, this may severely limit quasi-stars as a potential heavy seed formation mechanism – and thus an SMBH – formation pathway.
@@
Our models are presently a first-principles approximation of quasi-star evolution and black hole seed growth. We are limited by our understanding of accretion physics as well as the modeling of the interplay between mass ejection and accretion at the surface of the quasi-star. We plan to explore this interplay further in future work in order to understand the range of black hole seed masses that can be attained and the relevant conditions. We also plan to explore new inner boundary conditions as there is no indication that either of the two presented here is more accurate than any other. We have also neglected the roles of rotation and magnetic fields on quasi-star evolution, which we also aim to explore in future work.
@@
This analysis contained within this paper was presented as a poster at the Simons Foundation SCEECS Annual Meeting in February 2025 and published as a Master’s thesis in May 2025. Between then and the submission of this article we have learned of another group working on a similar problem whose work we would like to acknowledge here (J. Hassan et al., in prep.). Financial support for this publication comes from Cottrell Scholar Award #CS-CSA-2023-139 sponsored by Research Corporation for Science Advancement. This work was supported by a grant from the Simons Foundation (MP-SCMPS-00001470) to MC. This research was supported in part by the National Science Foundation under Grant No. NSF PHY-1748958. P.N. acknowledges support from the Gordon and Betty Moore Foundation and the John Templeton Foundation that fund the Black Hole Initiative (BHI) at Harvard University where she serves as one of the PIs.

### v3

@@ VI.4 Results
The results of implementing the Dutch wind scheme are plotted in Figs. 4 and 5. These winds are now strong enough to severely limit the final black hole masses. We see that the winds are enough to immediately eject the envelope upon expansion in the case of the $\alpha=1.4$ as it is enough to break the equilibrium that is otherwise reached. The $\alpha=1.0$ model lives only marginally longer, shedding its mass, while the black hole rapidly consumes what little remains. Our $\alpha<1.0$ models remain largely unaffected due to both their lower accretion luminosity and their increased temperature. As stated previously, the transition between wind models for the Dutch scheme occurs at $T_{\mathrm{eff}}=10^{4}$ K, which both low $\alpha$ models remain well above for the entirety of their lives. The Dutch scheme then serves as an intermediate case for a wind mass loss rate.
Figure 5: Kippenhahn diagrams showing the profile of the quasi-star using the saturated-convection boundary and including the Dutch wind scheme in both radial (left) and mass (right) coordinates. We show the Schwarzschild region (black), boundary region (grey), envelope (white), and the photosphere (black line) for models using $\alpha=0.8$ (top) and $\alpha=1.0$ (bottom).
@@
In addition to these global mass loss schemes, it is also possible for a local region of the envelope’s luminosity to exceed the Eddington limit and expel mass eruptively. This type of mass loss is supported by 3D radiation-hydrodynamic simulations and was also adapted to 1D for use within MESA by Jiang_2015; Jiang_2018 and Cheng_2024 respectively. A study done in parallel to ours by hassan explored this scheme further in the context of quasi-stars. They find a decrease in mass by a factor of $\sim 20$ in their equivalent of our $\alpha=1$ model. This sits within our range of wind schemes and represents a scenario between the Reimers and Dutch schemes. However, this process may be present alongside other wind drivers and become subdominant to the more severe mass loss rates of others.
@@
It is of note that envelope accretion can heavily suppress the outflow and sustain the quasi-star’s lifetime, allowing the black hole to accrete more of the envelope before it is shed (Dotan_2011; Fiacconi2015; Haemmerl__2021). This serves as motivation to pursue the addition of an envelope accretion scheme to our models in future work.
VII Discussion & Conclusions
@@
In this work, we have successfully implemented quasi-star models previously created by ball2012quasistars using the Cambridge STARS code (STARS) into MESA (Paxton et al., 2011). This implementation includes the Bondi inner boundary condition as well as the Reimers wind scheme, both of which serve as our fiducial case. We also add the ability to take the accretion rate as a semi-free parameter in order to proxy the addition of other physics. In our fiducial models, we show that final black hole masses of $\sim 10^{3}$ M can be attained even with accretion suppressing effects (e.g., angular momentum).
@@
We have implemented new inner boundary conditions based on a saturated-convection zone around the black hole proposed by coughlin2024. Doing so allows the BH to grow to $\sim$0.55 M, a factor of 5 increase from the previous $\sim$0.11 M achieved using the Bondi sphere as the inner boundary. The quasi-star, under these conditions, then becomes a heavy seed, capable of growing and reaching SMBH masses by early epochs thus explaining recent JWST observations. However, these results change dramatically when accounting for additional physics such as mass loss due to winds.
@@
We have studied three wind schemes in our models: the Reimers, Dutch, and super-Eddington radiation-driven winds. These serve as lower, middle, and upper bounds on mass loss rates for quasi-stars, as the actual values for surface mass loss are unknown without additional hydrodynamic and radiation transport simulations. We also presently do not account for envelope accretion, which is capable of balancing the mass loss and sustaining the envelope for an extended period of time. The Reimers and Dutch schemes may then serve as a proxy for that, given that the radiation-driven wind scheme is likely more realistic and catered to quasi-stars. We find that for the lower bound Reimers wind scheme, little change occurs to the final black hole mass for all accretion factors. In this case, the final SMBH masses will remain largely unaffected and are only limited by the size of the haloes that would would have formed the quasi-star in the first place. The Dutch scheme, however, severely limits the growth of our higher $\alpha$ values, suggesting that if net mass loss occurs in this way that accretion enhancing effects (e.g., magnetic fields and jets) may actually limit the viability of quasi-stars to form heavy seeds. If $\alpha$ remains less than one, the final black hole mass is decreased by a factor of $\sim 2-3$, and would result in a lower upper limit on SMBH masses at a given redshift by a similar factor if we assume all heavy seeds are formed through this mechanism. When $\alpha\gtrsim 1$, the scenario is similar to that of a light-seeding mechanism; however, the rapid winds would leave the stellar mass remnant in an insufficient environment to accrete rapidly enough to become an SMBH by the appropriate redshifts. Quasi-stars can reasonably span a range of $\alpha$ values, which would result in both a decrease in the number density as well as the upper mass limit for SMBHs at a given redshift. The exact effect on the SMBH mass function across redshifts requires deeper population analysis and is outside the scope of this work. Finally, the radiation-driven wind scheme presents the harshest limit on black hole growth, only allowing for a factor of two growth at most for all rates. If this wind is the most realistic, and accretion onto the envelope is not sufficient enough, this may severely limit quasi-stars as a potential heavy seed formation mechanism – and thus an SMBH – formation pathway.
@@
Our models are presently a first-principles approximation of quasi-star evolution and black hole seed growth. We are limited by our understanding of accretion physics as well as the modeling of the interplay between mass ejection and accretion at the surface of the quasi-star. We plan to explore this interplay further in future work in order to understand the range of black hole seed masses that can be attained and the relevant conditions. We also plan to explore new inner boundary conditions as there is no indication that either of the two presented here is more accurate than any other. We have also neglected the roles of rotation and magnetic fields on quasi-star evolution as they are non-trivial given both the 1D nature of MESA and the need to implement coupled schemes for the envelope and BH. We aim to explore the addition of these phenomena in future work.
@@
This analysis contained within this paper was presented as a poster at the Simons Foundation SCEECS Annual Meeting in February 2025 and published as a Master’s thesis in May 2025. Between then and the submission of this article we have learned of another group working on a similar problem whose work we would like to acknowledge here (hassan). Financial support for this publication comes from Cottrell Scholar Award #CS-CSA-2023-139 sponsored by Research Corporation for Science Advancement. This work was supported by a grant from the Simons Foundation (MP-SCMPS-00001470) to M.C. This research was supported in part by the National Science Foundation under Grant No. NSF PHY-1748958. E.F. acknowledges support from the Yale Center for Astronomy and Astrophysics Prize Fellowship. P.N. acknowledges support from the Gordon and Betty Moore Foundation and the John Templeton Foundation that fund the Black Hole Initiative (BHI) at Harvard University where she serves as one of the PIs.